THE LOW-LYING STATES OF SF_{n} SPECIES (n=1–6): INSIGHTS INTO HYPERVALENCY FROM THE RECOUPLED PAIR BONDING MODEL

\underline{D. E. WOON} and T. H. DUNNING, JR., Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

High level MRCI and RCCSD(T) calculations with correlation consistent basis sets were used to characterize SF_{n} species. By examining both the stable structures and the bonding processes that occur during SF_{n} + F → SF_{n+1} additions, we have derived a new model for describing hypervalent behavior that we call recoupled pair bonding, in which a pair of electrons on S can be decoupled to allow formation of a bond with F. The new model accounts for the origin of hypervalency, the presence of low-lying excited states, and the structures and spectral properties of neutral and ionic SF_{n} species; it has more predictive capability than other models. For example, while SF and SF_{2} both have covalently bonded ground states, they each have low-lying excited states with at least one recoupled pair bond. To the best of our knowledge, the SF(^{4}Σ^{−}), SF_{2}(^{3}B_{1}), and SF_{2}(^{3}A_{2}) states have not yet been observed experimentally.